OpenStudy (anonymous):

simplify the expression (1+cot theta)(1- cot theta) - csc^2 theta

OpenStudy (anonymous):

Multiple (1 - cot theta)(1+ cot theta) = 1 - cot^2 (theta). sue pythagorean identity that 1+ cot^2(theta) = csc^2 (theta). and you'll be done.

OpenStudy (anonymous):

use*

OpenStudy (mertsj):

\[(1+\cot \theta)(1-\cot \theta)-\csc ^2\theta =1-\cot ^2\theta-\csc ^2\theta =\] \[1-\cot ^2\theta-(1+\cot ^2\theta)=1-\cot ^2\theta-1-\cot ^2\theta =-2\cot ^2\]

OpenStudy (anonymous):

Awesome thank you guys!

OpenStudy (mertsj):

yw